We could not find problems with any individual piece of the REFL55 electronics chain, from photodiode to ADC. Nevertheless, the PRMI fringes witnessed by REFL55 is ~x10 higher than ~two weeks ago, when the PRMI could be repeatably and reliably locked using REFL55 signals (ETMs misaligned).
Discussion and next steps:
Q: Koji asked me what is the problem with this apparent increased optical gain - can't we just compensate by decreasing the whitening gain?
A: I am unable to transition control of the PRMI (no ETMs) from 3f to 1f, even after reducing the whitening gain on the REFL55 channels to prevent the saturation. So I think we need to get to the bottom of whatever the problem is here.
Q: Why do we need to transfer the control of the vertex to the 1f signals at all?
A: I haven't got a plot in the elog, but from when I had the PRFPMI locked last year, the DARM noise between 100-1kHz had high coherence with the MICH control signal. I tried some feedforward to try and cancel it but never got anywhere. It isn't a quantitative statement but the 1f signals are expected to be cleaner?
Koji pointed out that the MICH signal is visible in the REFL55 channels even when the PRM is misaligned, so I'm gonna look back at the trend data to see if I can identify when this apparent increase in the signal levels occurred and if I can identify some event in the lab that caused it. We also discussed using the ratio of MICH signals in REFL and AS to better estimate the losses in the REFL path - the Faraday losses in particular are a total unknown, but in the AS path, there is less uncertainty since we know the SRM transmission quite precisely, and I guess the 6 output steering mirrors can be assumed to be R=99%.